a particle experiences constant acceleration for 6seconds after starting from rest.if it travels a distance S1 in first 2seconds and a distance S2 in the next 2seconds and a distance S3 in the last 2seconds the S1:S2:S3

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

S1 = 1/3 S2 = 1/2 S3

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi